OpenStudy (anonymous):

I need to find an equation for the tangent to the curve at the given point. y=4-x^2 (-1,3)

OpenStudy (anonymous):

y' = -2x y'(-1) = -2*(-1) = 2 is your slope... so the eqn of tangent line is : y - 3 = 2(x+1)